Zaphod Beeblebrox has recently purchased a second-hand ballpoint pen business. In the first month, they purchased the following pens.

Color Count
Black 672
Blue 791
Red 163
Green 100
UltraViolet 258
Purple 101

What is the relative frequency of green pens purchased?

Answer in decimal form. Round to 3 decimal places as needed.
Answer #1

Solution :

Total = 672 + 791 + 163 + 100 + 258 + 101 = 2085

The relative frequency of green pens purchased is = 100 / 2085 = 0.048